OpenStudy (anonymous):

Find a possible solution to the equation sin(3x+13) = cos(4x). fan and metal

OpenStudy (anonymous):

good luck with that program is this a real question?

OpenStudy (anonymous):

i mean it cannot really be done without some sort of technology here are the answers http://www.wolframalpha.com/input/?i=sin%283x%2B13%29%3Dcos%284x%29

OpenStudy (anonymous):

i wasnt able to find an answer

zepdrix (zepdrix):

\[\Large\rm \sin(3x+13)=\color{orangered}{\cos(4x)}\]Since sine and cosine are co-functions, can't we just do this? \[\Large\rm \sin(3x+13)=\color{orangered}{sin\left(\frac{\pi}{2}-4x\right)}\]And then,\[\Large\rm 3x+13=\frac{\pi}{2}-4x\]Which gives us one solution... x=-1.6327...